OpenStudy (anonymous):

Use separation of variables to solve the initial value problem: dy/dx=(5-x^2)/(3y+5) and y=-1 when x=0 I'm not sure how to use separation of variables

ganeshie8 (ganeshie8):

cross multiply

ganeshie8 (ganeshie8):

\(\large \frac{dy}{dx} = \frac{5-x^2}{3y+5}\) \(\large (3y+5 )dy = (5-x^2)dx\)

ganeshie8 (ganeshie8):

integrate now

ganeshie8 (ganeshie8):

\(\large \frac{dy}{dx} = \frac{5-x^2}{3y+5}\) \(\large (3y+5 )dy = (5-x^2)dx\) \(\large \int (3y+5 )dy = \int (5-x^2)dx\)

OpenStudy (anonymous):

Thanks! I get the separation of variable now, I have one more question, when I integrate, I end with : (3y^2)/2 +5y=5x-(x^3/5) and I'm not sure how to solve for y

ganeshie8 (ganeshie8):

before solving "y", find the integration constant C, by using the initial conditions

ganeshie8 (ganeshie8):

when u do integration, the constant always pops up

ganeshie8 (ganeshie8):

you should get : (3y^2)/2 +5y=5x-(x^3/3) + C

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

when I solve for (0,-1) I get 3(-1)/2+5(-1)=0+c c=(-7/2) *I guess my question with this is if I'm allowed to just input the c on one side? and if my method works

ganeshie8 (ganeshie8):

correct !

ganeshie8 (ganeshie8):

(3y^2)/2 +5y=5x-(x^3/3) - 7/2

ganeshie8 (ganeshie8):

next, solve "y" if u want

ganeshie8 (ganeshie8):

or u may leave the solution in implicit form like above

OpenStudy (tkhunny):

1) Don't ever "cross multiply". Multiply bu things that are non zero and switch to the differential definition of the derivative. This gives: \(y \ne -5/3\) and then the integral expression you have created. 2) You seem to be struggling with your algebra. Try all that, again. \(\dfrac{3}{2}y^{2} + 5y = 5x - \dfrac{1}{3}x^{3} + C\) -- This is the expected result. 3) Use the initial values given to find C. \(\dfrac{3}{2}(-1)^{2} + 5(-1) = 5(0) - \dfrac{1}{3}(0)^{3} + C\). This is find, since we lose nothing by putting \(C\) on one side or \(C_{1}\) and \(C_{2}\) on either side. This suggests \(3/2 - 5 = C = -7/2\) Giving: \(\dfrac{3}{2}y^{2} + 5y = 5x - \dfrac{1}{3}x^{3} - \dfrac{7}{2}\) 4) Who said you had to solve for \(y\)? If you REALLY have to, I'm sure you've solved a Quadratic Equation, before. Give it a go.

ganeshie8 (ganeshie8):

depends on wat ur professor wants

OpenStudy (anonymous):

awesome! thanks so much! and I think I can leave it in implicit form, at least I hope

ganeshie8 (ganeshie8):

np :) if u r forced to solve, just use the quadratic formula